Fermi National Laboratory


Accelerator Update for August 22, 2005 - August 24, 2005

Monday, August 22
The day shift began with Operations monitoring store 4346 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using colliding beam, and with MiniBooNE, NuMI, and the Meson MCenter experiments taking beam.

Booster RF experts reconnected BRF4 at 8:53 AM.

Operations put the Recycler into shot setup to stash antiprotons at 12:11 PM.

The Recycler stash ended, Pbar resumed stacking, and MCenter resumed taking beam at 1:42 PM.

The Recycler began stashing antiprotons at 5:43 PM and ended at 6:23 PM. Pbar and MCenter resumed normal operations.

At 8:18 PM, Operations reported that a Linac RF station (KRF1) had tripped off for the first of many times over the next couple of hours on a LCW flow fault, but then eventually it settled down.

Tuesday, August 23
The midnight shift began with Operations monitoring store 4346 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 12:21 AM, KRF2 tripped off on a water flow fault for the first of six times over the next four hours. It reset each time, but not immediately.

At 2:18 AM, MCenter's analyzing magnets tripped off. It appeared to be a LCW problem and investigations began.

A water expert joined the MCenter leak investigation at 5:31 AM.

The water expert made up for the lost water in the Meson LCW system and MCenter resumed taking data at 6:07 AM. The leak investigation will continue later.

Pbar experts began a Debuncher study at 6:14 AM. This study interfered with stacking.

A Tevatron (TeV) expert began an end-of-store study at 9:07 AM.

The TeV study ended and Operations terminated store 4346 at 10:03 AM.

Operations put the TeV into shot setup at 10:15 AM.

Since MCenter was off due to shot setup, water experts began, at 10:23 AM, another search of MCenter for a possible LCW leak.

Operations established combination store 4347 at 1:05 PM with an initial luminosity of 120.81E30. MCenter and Pbar resumed normal operations.

A Recycler Electron Cooling study, which had started earlier, completed at 5:20 PM.

The Recycler began stashing antiprotons at 5:39 PM.

At 6:03 PM, a Pbar Debuncher RF station (DRF1) tripped off. Operators investigated and found a bad 10k-volt power supply. Stacking resumed at 6:38 PM. In the meantime, operators replaced the supply, but the new supply wouldn't power up. An expert suggested replacing the 20k-volt power supply, also. Operators did this and the station came back on line. Stacking continued through this repair.

The Recycler prepared to stash antiprotons at 11:52 PM.

Wednesday, August 24
The midnight shift began with Operations monitoring store 4347, with Pbar and the Recycler stashing antiprotons, and with MiniBooNE and NuMI taking beam.

At 12:39 AM, LRF2 tripped off for first of three times in an hour. The station reset each time, but not immediately.

The Recycler stash ended at 1:03 AM. Pbar and MCenter had resumed normal operations by 1:13 AM.

At 4:04 AM, the Tevatron Electron Lens tripped off. Operators investigated and discovered it had tripped on an airflow bit. It reset and was back on by 4:11 AM.

The Plans for Wednesday
The plans for today are to store, stack, and stash. There is a good chance that the Run Coordinator will allow an eight hour Pbar access on Thursday (8/25/05). This would mean no stacking for eight hours and no beam to NuMI, MiniBooNE, and MCenter for four to five hours (due to the access's proximity to the 8 GeV line).

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/24/2005   email Fermilab